After one interest period, the interest earned on a $7000 investment exceeds the interest earned on a $5000 investment by $160. The interest rate for the $5000 investment is 1.6% greater than the interest rate for the $5000 investment. Find the interest rates for the two investments.
Answer by lwsshak3(11628)

After one interest period, the interest earned on a $7000 investment exceeds the interest earned on a $5000 investment by $160. The interest rate for the $5000 investment is 1.6% greater than the interest rate for the $7000 investment. Find the interest rates for the two investments.
***
let x=interest rate for the $7000 investment
x+.016=interest rate for the $5000 investment
7000x-5000(x+.016)=160
7000x-5000x+80=160
2000x=80
x=80/2000
x=.04
x+.016=0.056
interest rate for the $7000 investment=4.0%
interest rate for the $5000 investment=5.6%